OpenStudy (anonymous):

Solve the system by substitution: -4.5x-2y=-12.5 3.25x-y=-0.75

OpenStudy (anonymous):

solve 3.25x-y=-0.75 for y

OpenStudy (anonymous):

okay would that be?... @♂ 3.25x-y=-0.75 -3.25x -3.25x -y=-0.75-3.25x

OpenStudy (anonymous):

\(-4.5x-2y=-12.5 \\3.25x-y=-0.75\) \(3.25x+0.75=y\) now you can plug that into the first equation

OpenStudy (anonymous):

Like this? @♂ -4.5x-2(3.25x+0.75)=12.5

OpenStudy (anonymous):

exactly except the 12.5 is negative

OpenStudy (anonymous):

Okay cool thank you! Is that all that I have to do for the answer? Or is there more steps? @♂

OpenStudy (anonymous):

well you have to solve for y and x

OpenStudy (anonymous):

How exactly would I do that? I'm sorry I am not the best at this... @♂

OpenStudy (anonymous):

I messed up my math, I'ms upposed to have x=1

OpenStudy (anonymous):